A 0.1358-g portion of potassium iodate (MW 214.00), about 2 g of potassium iodide, and 2 mL of 6 M hydrochloric acid were dissolved in 25 mL of distilled water. The triiodide formed during the ensuring reaction was titrated to the starch endpoint with 31.94 mL of a thiosulfate solution. A 25.00-mL triiodide sample solution was titrated to the endpoint with 21.33 mL of the standardized thiosulfate solution. Calculate the concentrations of the thiosulfate solution and the triiodide solution.
Answer #1

Since, KIO3 contain 59.326% I-.

In, 2g of KIO3 contain 1.18652 g of I-.

Since, KI contain 76.49% I-.

In, 0.1358g of KI contain 0.1039 g of I-.

Initial I- concentration = 1.18652 + 0.1039 = 1.29042 g of I3-

1.29042g at I3- can be titrated using 31.94 mL thiosulfate solution

25 mL distilled water and 2 mL HCl = 27 mL = 0.027 L

[I3-] =1.29042 g / (380.7 g/mole x 0.027 L)

[I3-] = 0.12554 mole/L = 0.12554 M

27 mL I3- 0.12554 M solution can be titrated using 31.94 mL Na2S2O3 (X) M solution.

VNa2S2O3 = 21.33 mL

VI3- = 25 mL

CNa2S2O3 x VNa2S2O3 = CI3- x VI3-

CNa2S2O3 x 21.33 mL = 0.12554 M x 25 mL

CNa2S2O3= (0.12554 M x 25 mL) / 21.33 mL

CNa2S2O3 = 0.14714 M

ANSWERS:

[Na2S2O3] = 0.14714 mole/L

[I3-] = 0.12554 mole/L